19) The following information was provided by Jimbob Co. for the year ended.

COG manufactured 500,000

Ending finished goods inventory 100,000

Sales 800,000

Gross margin 200,000

What was the beginning finished goods inventory?

  1. 300,000
  2. 200,000
  3. 100,000
Answer #1

Gross Margin = Sales - Cost of Goods Sold

200,000 = 800,000 - Cost of Goods Sold

Cost of Goods Sold = 800,000-200,000 = 600,000

Ending finished goods inventory = Beginning finished goods inventory + Cost of Goods manufactured - Cost of Goods Sold

100,000 = Beginning finished goods inventory + 500,000 - 600,000

Beginning finished goods inventory = 100,000 - 500,000 + 600,000

= 200,000